OpenStudy (anonymous):

need the homogeneous solution (t^2)y''-ty'=ln(t)

OpenStudy (anonymous):

pleaseeeeee

OpenStudy (lalaly):

r(r-1)-r=0 r^2 - r-r =0 r^2-2r=0 r(r-2)=0 r=0 r=2 y=c1 + c2 t^2

OpenStudy (anonymous):

good stuff thanks

OpenStudy (lalaly):

your welcome :)